OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y - 4 = 3(x - 2) y - 2 = 3(x - 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

OpenStudy (anonymous):

Y-4=3(x-2)

OpenStudy (anonymous):

Thank you, can you explain to me how you got that?

OpenStudy (zale101):

do u know what formula @Katherine97 used>

OpenStudy (anonymous):

Y= Mx + B I think

OpenStudy (zale101):

no

OpenStudy (zale101):

y-y1=m(x-x1) what's this formula called?

OpenStudy (zale101):

m= slope (x1,y1) =point given

OpenStudy (anonymous):

I have no clue what that is called :( Could you explain it to me?

OpenStudy (zale101):

It is called "point-slope form".

OpenStudy (anonymous):

Yes :) If I need help once more, may I tag you, math is difficult for me, but I try lol

OpenStudy (zale101):

the answer should be like this y-4=3(x-2) y-4=3x-6 distribute y=3x-6+4 y=3x-2

OpenStudy (zale101):

ok, sure u can :-)